Night Report 2016-07-31

Observing: Morgan Fraser, Tom Reynolds

Support Team: Seppo Mattila

TAT: Lluís Galbany, Stefan Taubenberger

Weather conditions

Start of night: Thin cirrocumulus covering most of sky, seeing ~1"

Middle of night: Closed due to clouds

End of night: Clear, seeing 1-1.5

------------------------------------------------

Calibrations

Evening: EFOSC biases, spec flats

Software error caused 5.0" slit to be used instead of 1" throughout, so only 1.5" data is good

Clouds and this problem prevented evening skyflats

------------------------------------------------

Start of night: Had to restart observing system multiple times and wait for TOś to try and solve 1" slit problem.

No time for evening std, twilight taken up by attempt to fix system.

COULD NOT SOLVE PROBLEM: all spectra taken with 1.5" slit, 1.0" slit unavailable

23:00 SN2016ase (follow-up): Could not acquire even with deeper acquisition image due to clouds

23:15 Image analysis

23:30 Standard, EG274, 1.5" slit only

23:33 End of twilight

23:40 SN2016AJ (follow-up): Could not see target in a long R acquisition image, so did not continue. Assume we need cloudless night

23:47 ASASSN15oz (follow-up): 3x200s R, can see the target.

00:02 Gaia16awz (Classification): 300s, Gr13, 1.5" slit. Not enough signal, took 1200s same setup. Still could not extract spectrum

00:35 Gaia16awy (Classification): 1200, Gr13, 1.5" slit. CV, spectrum observed through clouds

01:02 Gaia16awm (Classification): 1200 Gr13, 1.5" slit, CV

01:32 AT2016ekg (Classification): 300s Gr13, 1.5" slit, cut short by closure. Early Ia, few days pre-max.

01:43 CLOSED DUE TO CLOUDS

04:05 REOPENED

04:05 SN2009ip (follow-up): Could not see the galaxy clearly in acquisition, aborted due to cloud

04:14 OGLE-131 (follow-up): Again aborted due to cloud

04:18 PS16dnm (classification): Could not acquire, even in a deep 60s exposure. Didn't look like clouds, target seems faint

Clouds almost all dispersed

04:24 PS16dnp (classification): 600s, Gr13, 1.5" slit. Early Ia, few days pre-max.

04:42 PS16dno (classification): Could not see target in acquisition image, presumably faded.

04:48 SN2009ip (follow-up): Observed 10x150 B, 8x150 V

05:48 OGLE-131 (follow-up): Observed 9x200 B

06:28 SN2016AJM (follow-up): Observed 4x300 B

06:54 ASASSN15og (follow-up): Observed BVgri, Gr16 2700s 1.5" x 2. Did not get spectroscopic flats, multiple problems with saturation and arc lamps being on

10:00 OGLE15xl (follow-up): Observed z 200sx9

10:00 Beginning of twilight